### v1

@@ 2.1 Setup
We assume $\Sigma(\mathcal{E})$ is known, which is standard in experimental planning (Gerber and Green, 2012) (in practice, any consistent estimator, e.g., from pilot studies would suffice). In addition, we assume that $\Sigma(\mathcal{E})$ is uniformly bounded, which implies that once we commit to learn a set of experimental parameters $\tilde{\theta}_{\mathcal{E}}^{\mathrm{exp}}$, their variance is finite. In addition $\Sigma$ is strictly positive definite, therefore assuming that the variance is bounded away from zero. [^6] For this latter condition to hold in an asymptotic framework with growing sample size, $\theta$ and $\tilde{\theta}$ can be defined as parameters of interest after appropriately rescaling by the square-root of the sample size; in this case $\Sigma$ denotes the asymptotic variance. See Section 4.1 for details.
We let $\mathcal{E}\in\mathcal{S},\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})$ for some constraint sets $(\mathcal{S},\mathcal{G}(\mathcal{E}))$ encoding feasibility or budget constraints. In our framework, restrictions on the experiments researchers can run $\mathcal{E}\in\mathcal{S}$ (and on their precision $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})$) can take any desired form. In most applications, we think of such constraints as arising from fixed costs of running an experiment and/or standard lower bound constraints on power for experimental estimates, which are common in empirical practice (Duflo et al., 2007; Athey and Imbens, 2017; List et al., 2011)
We write compactly $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$ indicating $\mathcal{D}$ the set of feasible experiments (e.g., sample size, etc.), with $\gamma\in\mathbb{R}^{|\mathcal{E}|}$ for a choice of experiments $\mathcal{E}$.
Our goal is to optimize both over the design and the weights $\gamma$. Before introducing our main research question, we impose the following assumption.
Assumption 1 (First-order estimation error).
For any admissible $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume
for known weights $\omega\in\mathbb{R}^{p}$, with $|\omega_{j}|>0$ for all $j$.
Assumption 1 holds exactly for linear $\tau(\cdot)$ and serves as a first-order approximation for smooth nonlinear $\tau(\cdot)$.
Remark 1 (Non-linear $\tau(\theta)$).
Section 4.1 (and Example 3 below) extends our framework to non-linear estimands and shows how Assumption 1 serves as an approximation via a Taylor expansion within a local asymptotic framework, where the bias is local to zero, but at a possibly faster, equal or even slower rate than the estimator’s standard error. For non-linear $\tau$, $\omega$ is replaced by the gradient of $\tau$ evaluated at the baseline observational estimates (i.e., $\omega=\frac{\partial\tau(\theta)}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}+o_{p}(1)$), which is typically observed before the experiment is conducted.
Remark 2 (When researchers can only identify linear combinations of the parameters).
In some applications, researchers may be able to identify via an experiment a linear combination of the parameters of interest (up-to a first order approximation), defined $\phi\equiv W^{\top}\theta$, where $W\in\mathbb{R}^{p\times s}$ is a known matrix. Suppose that $\widetilde{\phi}$ is an unbiased estimator of $\phi$, and the researcher can access a subset of its entries via an experiment.
Our framework extends to this setting as follows. Define the coefficient of the linear projection of $\omega$ onto the column space of $W$: $\pi=\left(W^{\top}W\right)^{-1}W^{\top}\omega$ (known ex-ante). Define the residual $v=\omega-W\pi$. Then,
$\tau=\left[W\pi+v\right]^{\top}\theta=\pi^{\top}\phi+v^{\top}\theta\implies\tau=\check{\omega}^{\top}\check{\theta},$
where $\check{\omega}\equiv(\beta^{\top},v^{\top})^{\top}$ and $\check{\theta}=(\phi^{\top},\theta^{\top})^{\top}$. This corresponds to our original setting where the researcher’s feasible set only allows access to unbiased estimates of a subset of the parameters, corresponding to the entries of $\phi$. ∎
2.2 Research questions
@@
    Preliminary step: weights. For each $(\mathcal{E},\Sigma(\mathcal{E}))\in\mathcal{D}$, choose $\gamma$ to combine experimental and observational estimates on the selected coordinates.
@@
    Middle step: precision. For each $\mathcal{E}$, choose $\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})$, where $\mathcal{G}(\mathcal{E})$ denotes an arbitrary set of positive definite covariances under feasibility constraints.
@@
    Outer step: experiment choice. Choose $\mathcal{E}\in\mathcal{S}$, the feasible set of parameters to learn experimentally for given arbitrary constraint set $\mathcal{S}$.
Choosing $\gamma$ (step 1) for a fixed design follows a long-standing tradition in econometrics and statistics (Armstrong et al., 2024; Andrews and Shapiro, 2021; Donoho, 1994; Athey et al., 2025). We study this question here in combination with the experimental design problem, step 2 and 3, which has not been studied by the references above; this, as we show, will change the underlying optimization problem (also for $\gamma$). The central challenge is that performance depends on the unknown bias $b$ in the observational estimates.
2.3 Mapping to a simple two parameters model
While we derive our results in general form, we will build intuition using a stylized description with a two-parameter model.
Setting 2 (Illustration with two-parameter model).
Consider two parameters $\theta=(\theta_{1},\theta_{2})^{\top}$, and mutually independent observational and experimental estimates
$\tilde{\theta}^{\mathrm{obs}}_{j}-\theta_{j}\sim\mathcal{N}(b_{j},\ \sigma_{j}^{2}),\qquad\tilde{\theta}^{\exp}_{j}-\theta_{j}\sim\mathcal{N}(0,\ v_{j}^{2}),\qquad j=1,2.$
We may run one experiment: $\mathcal{S}=\{\{1\},\{2\}\}$. If we pick $j$, we estimate
$\hat{\theta}_{j}=\gamma_{j}\,\tilde{\theta}^{\exp}_{j}+(1-\gamma_{j})\,\tilde{\theta}^{\mathrm{obs}}_{j},\qquad\hat{\theta}_{-j}=\tilde{\theta}^{\mathrm{obs}}_{-j},$
and, to first order, $\tau(\theta)-\tau(\hat{\theta})=\omega_{1}(\theta_{1}-\hat{\theta}_{1})+\omega_{2}(\theta_{2}-\hat{\theta}_{2}).$
@@ 2.3 Mapping to a simple two parameters model
Let $\tilde{\theta}^{\mathrm{obs}}=(\tilde{\theta}^{\mathrm{obs}}_{1},\tilde{\theta}^{\mathrm{obs}}_{2})^{\top}$ denote observational estimates obtained in Gechter et al. (2024) from a structural model and potentially biased due to misspecification: $\mathbb{E}[\tilde{\theta}^{\mathrm{obs}}]-\theta=(b_{1},b_{2})^{\top}.$ A budget constraint allows an experiment in only one site, so $\mathcal{S}=\{\{1\},\{2\}\}$. If site $j$ is chosen, we obtain an unbiased $\tilde{\theta}^{\mathrm{exp}}_{j}$; the other site $k\neq j$ remains observational. Given $\mathcal{E}\in\mathcal{S}$,
@@
(Bergquist and Dinerstein, 2020) conduct demand and supply experiments in food markets in Kenya. Suppose here we are interested in similar applications in Uganda. For exposition, consider a basic linear demand and supply
@@
Due to cost constraints, $\mathcal{S}=\{\{D\},\{S\}\},$ i.e., we can learn either demand (estimating $\tilde{\beta}_{D}^{\mathrm{exp}}$, with randomized price discounts) or supply (estimating $\tilde{\beta}_{S}^{\mathrm{exp}}$, by introducing regulatory cost shocks on firms). Let
$\big(\tilde{\theta}^{\mathrm{obs}}-\theta\big)\ \sim\ \mathcal{N}\big(b,\Sigma\big),\quad\tilde{\theta}^{\mathrm{obs}}\equiv\sqrt{n}(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top},\quad\theta\equiv\sqrt{n}(\beta_{D},\beta_{S})^{\top}$
with $\theta$ and $\tilde{\theta}$ denoting the parameter and estimator here rescaled by the square-root of the sample size, and $b$ capturing bias. Our goal is to choose whether to conduct a supply or demand experiment in Uganda to estimate $\tau$. For given estimators $\hat{\theta}$ that combine the estimates from Kenya with the chosen experimental estimate from Uganda, under local asymptotics and a first-order Taylor expansion described in Section 4.1, we can write [^7] Here, we are using a local asymptotic framework where the bias of the estimates in Kenya is non-negligible but small relative to $n$, i.e., $b/\sqrt{n}=o(1)$, so that $\omega$ can be consistently estimated using observational data.
$\tau-\widehat{\tau}=\underbrace{\frac{1}{\sqrt{n}}\check{\omega}(\mathrm{plim}(\tilde{\beta}^{\mathrm{obs}}))^{\top}(\theta-\widehat{\theta})}_{\text{main first order effect}}+\underbrace{o_{p}\left(\frac{b}{n^{1/2}}\right)}_{\text{small higher order effects}},\qquad\check{\omega}(\beta)=\frac{t}{(\beta_{D}+\beta_{S})^{2}}\begin{pmatrix}-\beta_{S}\\ \beta_{D}\end{pmatrix}$,
@@ 3 Robust experimental design
In this section we introduce an experimental design focusing on the mean-squared error (MSE) of the estimator $\hat{\tau}$, a common measure of precision. In Section 4.2 we extend the framework to minimize the length of the confidence intervals.
3.1 Adaptation regret for experimental design
For a given design $(\mathcal{E},\Sigma(\mathcal{E}))$ and shrinkage weights $\gamma$, define the MSE at a fixed observational-bias vector $b$ as
@@ 3.1 Adaptation regret for experimental design
where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$. The MSE is a measure of precision of $\hat{\tau}$ and ideally, one would minimize $\mathrm{MSE}_{b}$ over $(\mathcal{E},\Sigma(\mathcal{E}),\gamma)\in\mathcal{D}$.
However, because $b$ is unknown, we first consider an uncertainty set given by an $\ell_{\infty}$-ball,
@@
with $B\geq 0$ an upper bound on the largest coordinate-wise bias. The $\ell_{\infty}$ choice has a desirable property relative to $\ell_{1}$ or $\ell_{2}$: it does not force a trade-off across coordinates (a large bias in one component need not be “offset” by a small bias elsewhere), which is attractive when biases may be positively correlated across parameters. The drawback is that worst-case solutions depend on the radius $B$ which may be unknown in practice.
@@
However, having to specify $B$ can pose a large burden on the researchers and make the choice of the experiment sensitive to $B$. Nevertheless, we could seek designs that perform as close as possible to the oracle that knows $B$, uniformly over the values of $B$.
This idea of performing as close as possible to the oracle is based on an extensive literature on regret minimization (Manski, 2004; Manski and Tetenov, 2007; Montiel Olea et al., 2023; Kitagawa and Tetenov, 2018; Manski and Tetenov, 2016). We minimize the worst-case proportional increase in MSE relative to the oracle across all bias radii:
@@
Whereas the above references consider the choice of an estimator for fixed design, here, the regret is relative to an oracle that chooses both the estimator and the design. Our key contribution – different from the references above – is the study of the optimal experimental design informed by biased (observational) estimates. This leads to a different definition of regret optimization, both because we optimize over the class of designs and $\Sigma$, but also in $\gamma$ due to the different definition of oracle. The optimal solution consists of a pair of design and estimator that can be pre-specified.
3.2 Optimal experimental design: main result
Before stating the result, we introduce compact notation. For a given $(\mathcal{E},\gamma)$, with $\gamma\in\mathbb{R}^{|\mathcal{E}|}$ define
$\tilde{\omega}_{\text{obs},j}(\mathcal{E},\gamma)=\begin{cases}\omega_{j},&j\notin\mathcal{E},\\ \omega_{j}(1-\gamma_{j}),&j\in\mathcal{E},\end{cases}\qquad\tilde{\omega}_{\text{exp}}(\mathcal{E},\gamma)=\omega_{\mathcal{E}}\circ\gamma$
where $\circ$ denotes the component-wise product and $\omega_{\mathcal{E}}$ is the subvector of $\omega$ corresponding to the entries indexed by $\mathcal{E}$. Let
$\tilde{\omega}(\mathcal{E},\gamma)\equiv\begin{pmatrix}\tilde{\omega}_{\text{obs}}(\mathcal{E},\gamma)\\[2.0pt] \tilde{\omega}_{\text{exp}}(\mathcal{E},\gamma)\end{pmatrix}\in\mathbb{R}^{\,p+|\mathcal{E}|}.$ (3)
$\alpha(\mathcal{E},\Sigma,\gamma)\;\equiv\;\tilde{\omega}(\mathcal{E},\gamma)^{\top}\Sigma(\mathcal{E})\,\tilde{\omega}(\mathcal{E},\gamma),\qquad\beta(\mathcal{E},\gamma)\;\equiv\;\Big(\,\|\omega\|_{1}-\|\omega_{\,\mathcal{E}}\|_{1}+\,|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\,\Big)^{\!2},$ (4)
with $|1-\gamma|$ indicating the absolute value of each entry of the vector $1-\gamma$.
Define the best achievable variance and worst-case bias divided by $B^{2}$ as
$\alpha^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\ \min_{\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})}\ \min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\alpha(\mathcal{E},\Sigma,\gamma),\qquad\beta^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\ \min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\beta(\mathcal{E},\gamma)=\min_{\mathcal{E}\in\mathcal{S}}\Big(||\omega||_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2}.$ (5)
Here $\alpha^{\star}$ is the variance of the most precise feasible design, while $\beta^{\star}$ is the squared bias of the least biased design. [^8] The expression for $\beta^{\star}$ corresponds to the following experiment choice: choose $\mathcal{E}$ with the largest $\omega_{\mathcal{E}}||_{1}$ and set $\gamma_{\mathcal{E}}\equiv 1$ at those coordinates. Both quantities depend on the class of designs $\mathcal{S},\mathcal{G}(\mathcal{E})$ and are easy to compute: $\alpha^{\star}$ is a constrained quadratic minimization; $\beta^{\star}$ requires enumerating $\mathcal{E}\in\mathcal{S}$.
Ideally, we would like a design with variance equal to $\alpha^{\star}$ and bias equal to $\beta^{\star}$. Unfortunately, this may be infeasible as it might require different choices of experiments and estimators to achieve one or the other.
Instead, in our main theorem below we show that the regret-optimal pair of design and estimator trades-off the bias and variance and depends on how far each of these two components are from their smallest attainable value.

### v2

@@ 2 Problem description
We let $\mathcal{E},\Sigma(\mathcal{E})$ be within arbitrary constraint sets that may encode feasibility or budget constraints. In our framework, restrictions on the experiments researchers can run $\mathcal{E}$ and on their precision $\Sigma(\mathcal{E})$ can take any desired form. In most applications, we think of such constraints as arising from fixed costs of running an experiment and/or constraints on their power (Duflo et al., 2007; Athey and Imbens, 2017; List et al., 2011). We assume that $\Sigma(\mathcal{E})$ is uniformly bounded, which implies that once we commit to learn a set of experimental parameters $\tilde{\theta}_{\mathcal{E}}^{\mathrm{exp}}$, their variance is finite, and strictly positive definite, therefore assuming that the variance is bounded away from zero. [^5] For this latter condition to hold in an asymptotic framework with growing sample size, $\theta$ and $\tilde{\theta}$ can be defined without loss as parameters of interest after appropriately rescaling by the square-root of the sample size; in this case $\Sigma$ denotes the asymptotic variance. See Section 5.1 for details. The shrinkage estimator can be unrestricted (i.e., $\gamma$ can take any value on the real line) or restricted (when for example researchers aim to only use experimental evidence when available, setting $\gamma_{j}=1$ for $j\in\mathcal{E}$).
We extend our framework to more general methods of moments estimators in Section 4.
Remark 1 (Structure of $\Sigma$).
Because observational and experimental moments are typically computed from disjoint samples, in practice, it is natural to model the covariance as block diagonal, $\Sigma=\operatorname{diag}(\Sigma_{\text{obs}},\Sigma_{\text{exp}})$. The first block corresponds to observational estimates (fixed and not subject to design choices), while the second corresponds to experimental estimates, and it is a function of the sample size allocated to treatment arms. This is a special case of our framework which can be encoded in the constraint set $\mathcal{D}$. ∎
Research questions
@@
    Preliminary step: weights. For each feasible $\mathcal{E},\Sigma(\mathcal{E})$, choose $\gamma$ to combine experimental and observational estimates on the selected coordinates. (If researchers restrict their estimators to only use experimental evidence when available, set $\gamma=1$.)
@@
    Middle step: precision. For each feasible $\mathcal{E}$, choose $\Sigma(\mathcal{E})$ (i.e., sample allocation).
@@
    Outer step: experiment choice. Choose a feasible set of experiments $\mathcal{E}$.
Choosing $\gamma$ (step 1) for a fixed design follows a long-standing tradition in econometrics and statistics (Armstrong et al., 2024; Andrews and Shapiro, 2021; Donoho, 1994; Athey et al., 2025). We study this question here in combination with the experimental design problem, step 2 and 3; this, as we show, will change the underlying optimization problem (also for $\gamma$). The central challenge is that performance depends on the unknown bias $b$.
Mapping $\tau$
A key feature of our framework is that researchers explicitly parameterize $\theta$ and $\tau(\cdot)$, thereby pre-specifying and making transparent which biases drive the design problem; we see this as an advantage because it clarifies the role of the experiment in complementing existing evidence. The sensitivity of $\tau$ to $\theta$, captured by its gradient $\omega(\theta)=\partial\tau(\theta)/\partial\theta$, determines how the (first-order) bias propagates onto the final estimand. For simplicity, we impose exact linearity, $\tau(\theta)=\omega^{\top}\theta$; all results extend to smooth, nonlinear $\tau$ via a first-order expansion, as discussed below.
Assumption 2 (First-order estimation error).
For any $(\mathcal{E},\Sigma,\gamma)\in\mathcal{D}$, assume
for known weights $\omega\in\mathbb{R}^{p}$, with $|\omega_{j}|\in(0,\infty)$ for all $j$.
Assumption 2 holds exactly when $\tau(\cdot)$ is linear and serves as a first-order approximation for smooth non-linear $\tau(\cdot)$. Section 5.1 (and Example 3) extend the framework to non-linear estimands encompassing as special cases the local-misspecification setups in Andrews et al. (2020); Armstrong and Kolesár (2021). For non-linear $\tau$, we replace $\omega$ with the gradient of $\tau$ evaluated at a preliminary observational estimate, i.e., $\omega\;=\;\frac{\partial\tau(\theta)}{\partial\theta}\Big|_{\theta=\tilde{\theta}^{\mathrm{obs}}}\;+\;o_{p}(1),$ which is typically available before the experiment is conducted. That is, building on references above, $\omega$ captures the first-order effect of the observational bias $b$ on the bias of $\hat{\tau}$. [^6] To see how this assumption holds under a first-order approximation, define $\theta$ the parameter rescaled by $\sqrt{n}$ (and similarly its bias $b_{n}$, indexed by $n$), where $n^{-1/2}$ is the rate of convergence of the observational estimators. Then second-order effects are negligible (and Assumption 2 holds asymptotically) provided that $b_{n}^{2}/\sqrt{n}\rightarrow 0$. This implies that Assumption 2 holds when the asymptotic bias $b_{n}$ is proportional to, smaller or even much larger than the asymptotic variance, provided that $b_{n}$ does not grow at rate faster than $n^{1/4}$.
We conclude with three examples in simple two-parameter models.
@@
Let $\tilde{\theta}^{\mathrm{obs}}=(\tilde{\theta}^{\mathrm{obs}}_{1},\tilde{\theta}^{\mathrm{obs}}_{2})^{\top}$ denote observational estimates obtained in Gechter et al. (2024) from a structural model and potentially biased due to misspecification: $\mathbb{E}[\tilde{\theta}^{\mathrm{obs}}]-\theta=(b_{1},b_{2})^{\top}.$ A budget constraint allows an experiment in only one site, so that $\mathcal{E}\in\{\{1\},\{2\}\}$. If site $j$ is chosen, we obtain an unbiased $\tilde{\theta}^{\mathrm{exp}}_{j}$; the other site $k\neq j$ remains observational. Given $\mathcal{E}$,
@@
Bergquist and Dinerstein (2020) conduct demand and supply experiments in food markets in Kenya. Suppose here we are interested in similar applications in Uganda. For exposition, consider a basic linear demand and supply
@@
Due to cost constraints, $\mathcal{E}\in\{\{D\},\{S\}\},$ i.e., we can learn either demand (estimating $\tilde{\beta}_{D}^{\mathrm{exp}}$, with randomized price discounts) or supply (estimating $\tilde{\beta}_{S}^{\mathrm{exp}}$, by introducing regulatory cost shocks on firms). Let
$\big(\tilde{\theta}^{\mathrm{obs}}-\theta\big)\ \sim\ \mathcal{N}\big(b,\Sigma^{\mathrm{obs}}\big),\quad\tilde{\theta}^{\mathrm{obs}}\equiv\sqrt{n}(\tilde{\beta}_{D}^{\mathrm{obs}},\tilde{\beta}_{S}^{\mathrm{obs}})^{\top},\quad\theta\equiv\sqrt{n}(\beta_{D},\beta_{S})^{\top}$
with $\theta$ and $\tilde{\theta}$ denoting the parameter and estimator here rescaled by the square-root of the sample size, and $b$ capturing bias. Our goal is to choose whether to conduct a supply or demand experiment in Uganda to estimate $\tau$. For given estimators $\hat{\theta}$ that combine the estimates from Kenya with the chosen experimental estimate from Uganda under local asymptotics described in Section 5.1
$\tau-\widehat{\tau}=\underbrace{\frac{1}{\sqrt{n}}\check{\omega}(\mathrm{plim}(\tilde{\beta}^{\mathrm{obs}}))^{\top}(\theta-\widehat{\theta})}_{\text{main first order effect}}+\underbrace{o_{p}\left(\frac{1}{n^{1/2}}\right)}_{\text{small higher order effects}},\qquad\check{\omega}(\beta)=\frac{t}{(\beta_{D}+\beta_{S})^{2}}\begin{pmatrix}-\beta_{S}\\ \beta_{D}\end{pmatrix}$,
@@ 3 Robust experimental design
Next, we introduce an experimental design focusing on the mean-squared error (MSE) of the estimator $\hat{\tau}$, defined, for given bias level $b$ and design $(\mathcal{E},\Sigma(\mathcal{E}))$ as
@@
where $\mathbb{E}_{\mathcal{E},\Sigma,b}$ denotes expectation under the data-generating process implied by $(\mathcal{E},\Sigma(\mathcal{E}))$ and observational bias $b$. Section 5.2 shows that our results for MSE also extend when minimizing confidence intervals’ length.
Ideally, one would minimize $\mathrm{MSE}_{b}$. However, because $b$ is unknown, we consider an uncertainty set given by an $\ell_{\infty}$-ball,
@@
with $B\geq 0$ an upper bound on the largest coordinate-wise bias. In our framework, a key challenge is that biases may arise from multiple parameters and uncertainty sets defined by different norms may be possible. A key insight here is that the $\ell_{\infty}$ choice has a desirable property relative to $\ell_{1}$ or $\ell_{2}$: it does not force a trade-off across coordinates (a large bias in one component need not be “offset” by a small bias elsewhere), which is attractive when biases may be positively correlated across parameters. The drawback is that worst-case solutions depend on the radius $B$ which may be unknown in practice.
@@
However, having to specify $B$ can pose a large burden on the researchers and make the choice of the experiment sensitive to $B$. Therefore, we seek designs that perform as close as possible to the oracle that knows $B$, uniformly over the values of $B$, following a long-standing tradition in decision theory (e.g. Manski, 2004; Manski and Tetenov, 2007; Montiel Olea et al., 2023; Kitagawa and Tetenov, 2018). Specifically, we minimize the worst-case proportional regret
@@
There are two key distinctions relative to such references using the adaptation regret. First (and importantly) we optimize over the design itself rather than fixing it ex ante; consequently, regret is defined relative to an oracle that jointly chooses both the estimator and the design. Second, we allow biases to arise across multiple parameters (rather than a single one), which necessitates a careful specification of the uncertainty set; we adopt an $\ell_{\infty}$-ball, which, as we show yields transparent and computationally-attractive solutions.
3.1 Optimal design
A natural question is whether the adaptation regret provides natural trade-offs between the variance and the bias. To do so, denote $\mathbb{V}_{\mathcal{E},\Sigma}(\cdot)$ the variance operator for given $(\mathcal{E},\Sigma)$.
Definition 1 (Variance regret).
Let
$\alpha(\mathcal{E},\Sigma,\gamma)\;\equiv\;\mathbb{V}_{\mathcal{E},\Sigma}(\hat{\tau}_{\gamma}),\qquad\alpha^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\ \min_{\Sigma(\mathcal{E})\in\mathcal{G}(\mathcal{E})}\ \min_{\gamma\in\mathbb{R}^{|\mathcal{E}|}}\alpha(\mathcal{E},\Sigma,\gamma),$ (2)
with $\alpha^{\star}$ denoting the smallest feasible variance. We will refer to $\alpha/\alpha^{\star}$ as the variance regret.
The variance regret denotes the ratio between the variance of a given design and estimator relative to the smallest achievable variance. We provide a similar definition for the bias. Let $|1-\gamma|$ define the absolute value of each entry of the vector $1-\gamma$.
Definition 2 (Bias regret).
Define
$\qquad\beta(\mathcal{E},\gamma)\;\equiv\;\Big(\,\|\omega\|_{1}-\|\omega_{\,\mathcal{E}}\|_{1}+\,|1-\gamma|^{\top}|\omega_{\mathcal{E}}|\,\Big)^{\!2},\quad\beta^{\star}\;\equiv\;\min_{\mathcal{E}\in\mathcal{S}}\Big(||\omega||_{1}-\|\omega_{\mathcal{E}}\|_{1}\Big)^{2},$
with $\beta/\beta^{\star}$ defined as the bias regret.
[…]
An ideal design, would set the variance equal to $\alpha^{\star}$ and bias equal to $\beta^{\star}$. Unfortunately, this may be infeasible as it might require different choices of experiments and estimators to achieve one or the other. This raises the question of how to trade-off these two components.
This trade-off is formally characterized in our main theorem below.